# Zhang, Liu, and Zhao

### v3

@@ 11.2 Convergence of Enhanced SMM to B-stationarity
    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (5) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (5).
@@
    With $\varepsilon>0$, under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$. In particular, if $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are piecewise affine and $\Theta$ is a polytope, the same convergence rate holds without the two assumptions.
@@
12 Additional Experimental Results
For the unconstrained Newsvendor problem in Section 6.1, we set $(c_{b},c_{h})=(5,5)$ and consider another demand model
$\overline{Y}^{\prime}(X)=4\sin(\pi X_{1})+\max\{16X_{2},-20X_{2}\}+10,$
where the sine function can be interpreted as the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents another nonlinear trend with respect to $X_{2}$. In practice, we may distinguish seasonal features from the others and build the improved DR represented as the sum of multiple DRs that are constructed for separate groups of features. Thus, we employ the sum of two PADR(2,2) models and the sum of two RKHS-DRs as the improved DRs in this instance and denote them by PADR(2,2)+ and RKHS-DR+, respectively.
a. Basic Results. In Figure 1p, we report results with the training sample size $n$ increasing from 50 to 1000. All PADR and RKHS-DR methods except PADR(2,2) converge close to SIMOPT. Figure 1q shows the results of running PADR(4,4) and PADR(2,2)+ for 30 rounds each, where we report the mean cost of the top 15 with standard deviation for each trial. It can be seen that although PADR(4,4) has a stronger approximation ability, PADR(2,2)+, designed with prior information, obtains lower and stabler costs.
(p) Comparison with Benchmarks
(q) PADR(4,4) vs. PADR(2,2)+
Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$: Basic Results
b. Varying Feature Dimensions. We set sparse and dense models in the same way as in Section 6.1 and report the comparison results with RKHS-DR and RKHS-DR+ in Figure 1s. PADR(2,2)+ outperforms all the others in both cases. It is worth mentioning that PADR(2,2) is close to PADR(2,2)+ in high-dimensional cases, showing that the training difficulty gradually dominates compared to the value of prior information as the dimension grows. But RKHS-DR+ consistently performs better than RKHS-DR, and it performs better than PADR(2,2) when $p$ is small. As for the running time, PADR(4,4) and PADR(2,2)+ grow linearly with $p$ and are longer than RKHS-DR and RKHS-DR+ only when $p\geq 30$.
(r) Sparse Demand Model
(s) Dense Demand Model
Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$ and Varying Feature Dimensions

### v4

@@ 9.2 Convergence of ESMM to B-stationarity
    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (9.2) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (9.2).
@@
    With $\varepsilon>0$, under Assumptions (B5) and (B6), $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$. In particular, if $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are piecewise affine and $\Theta$ is a polytope, the same convergence rate holds without the two assumptions.
@@
For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (9.2) with probability 1 according to Theorem 9.8, and thus is a global minimizer of (9.2) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed. Furthermore, such a global minimizer of the penalized problem (9.2) should be also a global minimizer of (9.2) since it is feasible according to Proposition 9.6. Under such settings, the asymptotic consistency of the PADR framework for the constrained case in Theorem 9.3 turns into asymptotic optimality.
9.3 Experiments on constrained newsvendor problem
We compare PADR and other benchmarks on two-product constrained Newsvendor problems, assuming that the uncensored historical demand data is available. We denote the unit back-order cost, unit holding cost, and order decision by $c_{bi}$, $c_{hi}$, and $z_{i}$, respectively, for product $i=1,2$. The demand models for the two products are set as follows:
$\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30.$
In this subsection, we compare PADR with prescriptive methods and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
9.3.1 Newsvendor problem with a linear constraint.
We first consider an inventory capacity constraint for the two products, which is formulated as $z_{1}+z_{2}\leq C_{0}$. In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, such that the probability that the constraint is active at the optimal decision is around 0.5.
Since DR-based methods may yield infeasible decisions, in Figure 1u we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of the original decisions provided by PADR and RKHS-DR.
[…]
(s) Basic results
(t) Sparse demand model
(u) Dense demand model
Newsvendor problem with a linear constraint
With the basic demand model, PADR converges to SIMOPT and attains 100% feasibility when $n\geq 200$. In comparison, RKHS-DR and prescriptive methods display notable gaps to SIMOPT, with the feasible frequency of RKHS-DR being only around 80%. Figures 1t and 1u show the numerical results with varying feature dimensions. Similar to the unconstrained cases, PADR outperforms the other methods in the sparse case when $p<20$ and in the dense case when $p<50$. Notably, with penalization, PADR achieve 100% feasibility when $p<20$. Although PADR can not guarantee 100% feasibility of the decision compared to prescriptive methods, it still provides satisfactory performance via projection.
9.3.2 Newsvendor problem with a nonconvex constraint.
We next consider a capacity constraint $C(z_{1})+C(z_{2})\leq C_{0}$ with the concave capacity cost function (18) for the two products. We set $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and $C_{0}=50$ with the same probability for the constraint being active as the experiment in Section 9.3.1. Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1x, with a careful hyperparameter selection to ensure high feasible frequencies.
We use the standard MM algorithm to solve the penalized ERM problem of RKHS-DR, referred to here as RKHS-DR(MM). For prescriptive methods and SIMOPT, we solve the two-dimensional nonconvex-constrained problems using grid search. The sof method in Kallus and Mao (2023)is not suitable for this nonconvex-constrained problem, and it may require the study of the second-order perturbation analysis for stochastic optimization, which is beyond the scope of this paper.
For the problem with the basic demand model, both PADR and RKHS-DR attain 100% feasibility on the test set. PADR converges to SIMOPT with significantly fewer training samples than other benchmarks. Figures 1w and 1x show the results across varying feature dimensions. PADR outperforms the other methods in both sparse and dense cases when $p\leq 30$, achieving 90 % $\sim$ 100% feasibility in the sparse case when $p\leq 20$ and in the dense case when $p\leq 30$. Under the dense demand model, RKHS-DR performs well in low-dimensional cases, but its performance deteriorates as $p$ increases. Prescriptive methods—kNN (in the dense case), cart, and rf—maintain a stable yet noticeable gap to SIMOPT, which is possibly due to their limitations to approximate conditional distributions with the sample set that is not sufficiently large. The relatively small gap of PADR to SIMOPT shows PADR’s capability to learn the underlying decision rule structure in problems with nonconvex constraints while maintaining a high probability of feasibility when $p\leq 30$.
[…]
(v) Basic results
(w) Sparse demand model
(x) Dense demand model
Newsvendor problem with a nonconvex constraint
[…]